OpenStudy (anonymous):

use the distributive law to multiply: 3(12+5u+7v)

OpenStudy (anonymous):

36+15u+21v

OpenStudy (anonymous):

\[3(12+5u+7v)\] \[=36+15u+21v\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

3(12 + 5u + 7v)= (3*12) + (3 * 5u) + (3* 7v) = 36 + 15u + 21v